### v1

@@ 4.1 Test case 3: agglomeration capabilities assessment of test datasets
Figure 10: Test case 3: examples of agglomerated meshes, one for each dataset. The first one shown is always representing the physical parameter heterogeneities inside the domain and the others are the agglomerated elements grouped by main parameter value inside the element. Each figure is reported from two different views to better visualize the agglomerated elements.
In this test case, we consider three different test datasets, constructed on unitary cubes, but with different distribution of physical parameters. In particular, the datasets are composed as follows:
  • Dataset 1: $25$ tetrahedral meshes with a number of elements between $373$ and 10 3891038910\ 38910 389 with the same physical parameter in the whole domain.
  • Dataset 2: $25$ tetrahedral meshes with a number of elements between $437$ and 8 54985498\ 5498 549 with two different physical parameters in two regions of the domain.
  • Dataset 3: $50$ tetrahedral meshes with a number of elements between $529$ and 14 9601496014\ 96014 960 with two different physical parameters in more than two regions of the domain.
As an example, we report a random mesh for each of the presented datasets in Figure 10. Different colours are used to represent a different value of the physical parameter and the bold lines denote the region delimitations.
The results in this section are obtained using these datasets and clustering until reaching a diameter of the agglomerated elements of $0.25$ times the diameter of the original mesh $h$. In Figure 10, we report an example of mesh for each of the three test sets, together with the agglomerated meshes. The agglomerated meshes are visualized separating the different physical parameter regions, to underline the good level of separation of heterogeneities obtained by the GNN algorithm. In particular, only the example in Dataset 3 shows a single element containing two different physical values in the middle region. Finally, we can see that the agglomerated elements seem to have good shapes and to be quite uniform in all the considered cases.
@@
Table 5: Test case 3: average values of the computed quality metrics (CR, UF, VD, and HE) for the agglomerated grids reported in Figure 10.
To have a quantitative understanding of the agglomeration quality, we collect the average values of the quality metrics in the three different datasets in Table 5. First of all, the quantitative feedback confirms the good behavior of such a model in agglomerating while preserving the different physical regions. Indeed, on average, approximately only $2\%$ of the final elements is heterogeneous with respect to the physical parameters in the starting tetrahedral elements. This achievement is significant because it allows us to agglomerate the heterogeneous meshes with the awareness of obtaining a meaningful coarse mesh of the computational domain.
Moreover, the obtained values are coherent to what was obtained in the homogeneous case, with the values reported in Table 3. The important remark is that we do not make comparisons with the state-of-the-art methods because they are not able to perform automatically the agglomeration considering the heterogeneity in the physical parameter values. Therefore, the real strong point of this model is not the performance in terms of quality metrics, which can undoubtedly be improved by further developments, but the possibility of making this agglomeration process automatic, without a segregated agglomeration and a glue of the obtained meshes.
@@ 4.2 Test case 4: agglomeration of a mesh containing microstructures
Computational domains with complex microstructures at different scales are typical in the field of Integrated Computational Material Engineering (ICME). More specifically, many heterogeneous materials have this configuration, such as alloys, polymers, tissues, bones, and porous rocks [54, 55, 56, 57, 58].
The computational domain is a unitary cube with $15$ microstructures in random positions, meshed by a tetrahedral grid made by 71 8367183671\ 83671 836 tetrahedral elements (see Figure 11(a)). Each microstructure is composed of $10$ to $20$ tetrahedra, and so it is of negligible dimensions with respect to the diameter of the domain. We model this domain by setting the ”physical” parameter of the background to 00 and the microstructure one equal to $1$.
@@
Figure 11: Test case 4: domain with microstructure (a), and visual inspection of three domain’s regions containing a microstructure (b),(c), and (d).
We perform a mesh agglomeration process using the proposed GNN algorithm until reaching a diameter of the agglomerated mesh’s elements less than $0.15\,h$, where $h$ is the diameter of the original mesh. We get an agglomerated mesh of 2 26422642\,2642 264 elements. The first result to be underlined is the fact that our model, despite the microstructures having a small size, can identify each one of them. Indeed, where we have one of these particles in the initial grid, we found at least one agglomerated element with the physical parameter $1$ in the final mesh.
@@
In Figure 11, each shadowed color represents a different agglomerated element, the brighter colors are the agglomerates with physical parameter $1$, and, finally, the bold black lines represent the boundary of the microstructure. We observe that the model can gathers the tetrahedra of a microstructure in one agglomerated element as in Figure 11(b). In some cases, the elements of one structure are divided into smaller agglomerates (see Figure 11(c)). In some instances, we find some tetrahedra of a microstructure inserted in one of the adjacent background’s bigger elements as in Figure 11(d). However, by calculating the percentage of tetrahedra of each microstructure correctly classified with physical parameter $1$ in the final grid, we obtain a value around $84.90\%$.
4.3 Test case 5: agglomeration of a mesh of a human brain
@@
Figure 12: Test case 5: brain geometry with a distinction between white (in red) and grey (in blue) matters and agglomerated mesh. In the mesh visualization, the agglomerated elements are shown with different colors. We report a full brain visualization, a clipped view in the coronal plane, and one in the transverse plane. In the clipped views we also show separately the two matters agglomerated elements.
The final test proposed in this work is the application to a complex and heterogeneous domain. The brain is an intricated geometrical structure, with a natural distinction into two parts, namely, white and grey matter (see Figure 12). The computational domain can be reconstructed starting from MRI scans (in this work we consider one from the OASIS database [52]). The domain is also highly non-convex, and constructing efficient and high-quality tetrahedral meshes is complex. In particular, some applications require considering the heterogeneities in the cerebral medium for a correct physical description of the phenomenon [3, 59].
The computational domain is meshed by a tetrahedral grid made of 67 1646716467\,16467 164 elements, and the grid is constructed to preserve the distinction between white and grey matter. We model the domain by setting the grey matter tag to 00 and the white matter to $1$. We perform a mesh agglomeration process using the proposed GNN algorithm until reaching a diameter of the agglomerated mesh’s elements less than $0.15\,h$, where $h$ is the diameter of the original mesh. Finally, we get an agglomerated mesh of 3 23032303\,2303 230 elements.
In Figure 12, each shadowed color represents a different agglomerated element. We can observe a good distinction in the agglomeration with respect to the physical parameter describing the matter properties. A quantitative analysis confirms this quality. Indeed, by measuring the number of the tetrahedra agglomerated into an element with the wrong physical parameter, we can observe they are only the $2.0979\%$ of the total. Moreover, considering that almost the $35.4059\%$ of the initial mesh elements have a face on the interface between white and grey matters, we notice that only the $5.9253\%$ of the latter group is agglomerated incorrectly. These results are of particular significance, considering that none of the state-of-the-art methods can address the task automatically and that the brain geometry and the internal interface are highly complex.
@@ 5 Conclusions
In this work, we have developed a novel agglomeration algorithm for three-dimensional grids, via a graph partitioning algorithm based on Graph Neural Networks. The model is based on learning the geometrical information of the mesh’s elements automatically and cost-effectively, remaining independent of geometry and underlying differential models. This structure allows us to approach a wide range of situations, which would not be possible using classical strategies. Indeed, the proposed GNNs-enhanced algorithm can perform completely automatically the agglomeration of heterogeneous domains and ones containing microstructure, which is not possible by relying on classical methods. Another novelty of the proposed method consists of overcoming the main limitations of state-of-the-art approaches like METIS and k-means using GNN architectures, preserving the geometrical quality of the initial grid, and reducing the mesh complexity in terms of memory storage and associated computational costs.
First of all, we re-framed the problem of mesh agglomeration as a graph partitioning problem by exploiting the graph representation of the connectivity structure of the mesh’s elements. Then, we developed in this context a framework to employ GNN-based strategies that can exploit the geometrical information of the grid. After, we compared the performance of the proposed GNN strategies (properly trained) with k-means and METIS, classical algorithms for graph partitioning, over a set of different grids, featuring also complex real domains such as brain ventricles from MRI scans. Results show that the GNN proposed method can preserve a high mesh quality, with the advantage of processing naturally and simultaneously both the graph structure of the mesh and the geometrical information. Moreover, GNNs have a significantly lower online computational cost.
Finally, we refined the model to perform the agglomeration of a heterogeneous domain and tested it on different types of heterogeneous datasets, to confirm the quality of the constructed agglomerated meshes. Finally, we test the GNN algorithm on the micro-structures framework and a complex domain such as a real brain geometry. We remark also that the state-of-the-art methods cannot automatically perform the agglomeration of these domains.
Future developments certainly include improving the generalization capabilities of the GNNs, by generating a larger training dataset. Another possibility is to re-frame the approach in a reinforcement learning framework, as proposed by [60].

### v2

@@ 4.1 Test case 3: agglomeration capabilities assessment of test datasets
In this test case, we consider three different test datasets, constructed on unitary cubes, but with different distributions of physical parameters. In particular, the datasets are composed as follows:
  • Dataset 1: $25$ tetrahedral meshes with a number of elements between $373$ and 10 38910\ 389 with the same physical parameter in the whole domain.
  • Dataset 2: $25$ tetrahedral meshes with a number of elements between $437$ and 8 5498\ 549 with two different physical parameters in two regions of the domain.
  • Dataset 3: $50$ tetrahedral meshes with a number of elements between $529$ and 14 96014\ 960 with two different physical parameters in more than two regions of the domain.
Figure 10: Test case 3: examples of agglomerated meshes, one for each dataset. The first figure (top-left) in each block represents the physical parameter heterogeneities, and the others show the agglomerated elements. Each figure is reported from two different views to visualize the agglomerated elements better.
The results in this section are obtained using these datasets and clustering until reaching a diameter of the agglomerated elements of $0.25$ times the diameter of the original mesh $h$. In Figure 10, we report an example of a mesh for each of the three test sets, together with the agglomerated meshes. The first plot (top-left) in each block represents the initial mesh and the physical parameter heterogeneities, where different colors are used to represent different values of the physical parameter, and the bold lines denote the sub-region boundaries. The other plots in each panel show the agglomerated elements. Each figure is reported from two different views to have a good visualization of the agglomerated elements. The agglomerated meshes are visualized, separating the different physical parameter regions, to underline the level of separation of heterogeneities obtained by the GNN algorithm. In particular, only the example in Dataset 3 shows a single element containing two different physical values in the middle region. Finally, we can see that the agglomerated elements seem to be shape regular and to be relatively uniform in all the considered cases.
@@
Table 5: Test case 3: average values of the computed quality metrics (CR, UF, VD, and HE) for the agglomerated grids reported in Figure 10. The colors are scaled starting from the minimum value of the metric (red) to the maximum one (green).
To have a quantitative understanding of the agglomeration quality, we collect the average values of the quality metrics in the three different datasets in Table 5. First of all, the quantitative feedback confirms the good behavior of the proposed model in agglomerating while preserving the different physical regions. Indeed, on average, approximately only $2\%$ of the final elements are heterogeneous for the physical parameters in the starting tetrahedral elements. Moreover, the obtained values are coherent with what was obtained in the homogeneous case, with the values reported in Table 3. The important remark is that we do not make comparisons with either METIS or k-means here because they are not able to perform the agglomeration automatically considering the heterogeneity in the physical parameter values.
@@ 4.2 Test case 4: agglomeration of a mesh containing microstructures
Computational domains with complex microstructures at different scales are typical in the field of Integrated Computational Material Engineering (ICME). More specifically, many heterogeneous materials have this configuration, such as alloys, polymers, tissues, bones, and porous rocks [59, 60, 61, 62, 63]. The computational domain considered in this example is a unitary cube with $15$ microstructures in random positions, meshed by a tetrahedral grid made by 71 83671\ 836 tetrahedral elements (see Figure 11(a)). Each microstructure is composed of $10$ to $20$ tetrahedra, and so it is of negligible dimensions with respect to the diameter of the domain. We model this domain by setting the ”physical” parameter of the background to 0 and the microstructure one equal to $1$.
@@
Figure 11: Test case 4: domain with microstructure (a), and zoom of three domains’ regions containing a microstructure (b),(c), and (d).
@@
We agglomerate the underlying tetrahedral grid using the proposed algorithm until reaching a diameter of the agglomerated mesh’s elements less than 0.15h0.15\,h, where $h$ is the diameter of the original mesh. We get an agglomerated mesh of 2 2642\,264 elements. We observe that our model, despite the microstructures having a small size, can identify each one of them. Indeed, we found at least one agglomerated element per microstructure in the agglomerated mesh. In Figure 11, each shadowed color represents a different agglomerated element, the brighter colors are the agglomerates with physical parameter $1$, and, finally, the bold black lines represent the boundary of the microstructure. We observe that the model can gather the tetrahedra of a microstructure in one agglomerated element. In some cases, the elements of a microstructure are divided into smaller agglomerates (see Figure 11(c)). In some instances, a few tetrahedra of a microstructure are glued in one of the adjacent background’s bigger elements, as in Figure 11(d). However, by computing the percentage of tetrahedra of each microstructure correctly classified in the final grid, we obtain a value around $84.90\%$.
5 Robustness and generalization capabilities
In this section, we apply our algorithm to a diverse collection of meshes obtained from both heterogeneous and homogeneous domains to test its robustness and generalization capabilities.
5.1 Test case 5: mesh agglomeration of complex geometries
We consider computational surfaces obtained from different open-source datasets to test the ability of our method to generalize to various mesh sizes and geometrical shapes. We report all the information on the considered examples for this test case in Table 6. Moreover, we provide a visualization of the computational domains (and corresponding fine initial tetrahedral meshes) in Figure 12 (first row).
Domain Elements Database Source file type
$\Omega_{1}$ Arc de triomphe 179 582179\,582 FTetWild [64] Tetrahedral mesh (model 551021)
$\Omega_{2}$ Bolt 12 17512\,175 FTetWild [64] Tetrahedral mesh (model 371103)
$\Omega_{3}$ Damaliscus korrigum 507 480507\,480 Artec3D Surface PLY (meshed with VMTK [65])
$\Omega_{4}$ Human brain 609 770609\,770 OASIS-3 [57] MRI Image (meshed with SVMTK [66])
$\Omega_{5}$ Human rib cage 317 552317\,552 Free3D Surface OBJ (meshed with VMTK [65])
$\Omega_{6}$ Lucy angel statue 179 582179\,582 Stanford Repository Surface STL (meshed with VMTK [65])
$\Omega_{7}$ Nefertiti bust 351 086351\,086 Berlin Egyptian Museum Surface STL (meshed with VMTK [65])
Table 6: Test case 5: mesh information used in the test case and proper references to the databases.
Figure 12: Test case 5: computational meshes with indication of the heterogeneities (red and blue) for the domains $\Omega_{3}$$\Omega_{4}$, and $\Omega_{5}$ (first row) and corresponding agglomerated meshes (second row). Where the mesh edges are not shown, the surface is too refined to be visualized.
In this test case, we consider four homogeneous domains: the Arc de Triomphe ($\Omega_{1}$), a bolt ($\Omega_{2}$), the Lucy Angel statue ($\Omega_{6}$), and the Nefertiti bust ($\Omega_{7}$), and three heterogeneous ones
  • a damaliscus korrigum skull ($\Omega_{3}$), where we consider the difference between the bone and the horns;
  • a human brain ($\Omega_{4}$) with the distinction between white and grey matter, which is fundamental in the modeling of physical phenomena [3, 67];
  • a human rib cage ($\Omega_{5}$) with distinction of the thoracic vertebrae.
For all the geometries, we perform a mesh agglomeration process using the proposed GNN algorithm until reaching a diameter of the agglomerated mesh’s elements less than $0.15\,\textrm{diam}(\Omega)$, where $\textrm{diam}(\Omega)$ is the diameter of the domain. The agglomerated meshes are reported in the second row of Figure 12. In the visualization, each shadowed color represents a different agglomerated element. In Table 7, the number of elements of the agglomerated grids ($N_{\textrm{agg}}$) is compared with the corresponding number of elements of the initial tetrahedral meshes ($N_{\textrm{tets}}$). The last column of Table 7 also report the percentage reduction $\xi$ of the number of elements defined as
$\xi=\dfrac{N_{\mathrm{tets}}-N_{\mathrm{agg}}}{N_{\mathrm{tets}}}.$ (9)
Domain $N_{\textrm{agg}}$ $N_{\textrm{tets}}$ $\%$ reduction $\xi$
$\Omega_{1}$ Arc de triomphe 2 1602\,160 179 582179\,582 $98,7972\%$
$\Omega_{2}$ Bolt $170$ 12 17512\,175 $98,6037\%$
$\Omega_{3}$ Damaliscus korrigum 1 6831\,683 507 480507\,480 $99,6684\%$
$\Omega_{4}$ Human brain 22 52422\,524 609 770609\,770 $96,3061\%$
$\Omega_{5}$ Human rib cage 5 1925\,192 317 552317\,552 $98,3650\%$
$\Omega_{6}$ Lucy angel statue 1 6231\,623 179 582179\,582 $99,0962\%$
$\Omega_{7}$ Nefertiti bust $900$ 351 086351\,086 $99,7437\%$
Table 7: Test case 5: number of elements of the agglomerated grids ($N_{\textrm{agg}}$), corresponding number of elements of the initial tetrahedral grids ($N_{\textrm{tets}}$), and $\%$ reduction $\xi$ defined as in (9).
We observe that the agglomerates are coarser for homogeneous geometries with refined initial meshes and relatively simple boundaries (e.g., the Nefertiti bust). In contrast, the heterogeneous properties and complexity of the brain geometry result in a larger number of agglomerates. We remark that many of the meshes tested in this example are composed of more than $10$ times the elements of the most refined mesh in the training set ($\simeq 17\,000$ elements). In particular, the brain mesh is composed of more than $35$ times the elements of the finest mesh in the training set, confirming the scalability of the proposed method. This test case demonstrates the proposed algorithm’s robustness and generalization capabilities across a wide range of geometries and complexity levels. In the heterogeneous cases, we can also observe robustness concerning the physical parameters. To have a more quantitative analysis of the results of our GNN agglomeration strategy, we computed the number of tetrahedra on the initial mesh that were incorrectly agglomerated into an element with a different physical parameter. For the three considered cases, we have that this value is $0.0108\%$ for $\Omega_{3}$, $1.440\%$ for $\Omega_{4}$, and $0.1458\%$ for $\Omega_{5}$. The result of the brain geometry $\Omega_{4}$ seems to be worse. Still, we have to underline that almost $32.64\%$ of the initial mesh elements have a face on the interface between white and grey matter, leading to a particularly challenging interface to be considered. The good quality of the distinction between white and grey matter can be observed in Figure 13. These results are of particular significance, considering that neither the k-means nor METIS can automatically address this task and that the brain geometry and the internal interface are highly complex.
@@ 5.1 Test case 5: mesh agglomeration of complex geometries
Figure 13: Test case 5: brain geometry $\Omega_{4}$ with the corresponding distinction between white (in red) and grey (in blue) matters and corresponding agglomerated mesh. In the mesh visualization, the agglomerated elements are shown with different colors. We report full brain visualization, a clipped view in the coronal plane, and a transverse plane view. In the clipped views, we also show separately the agglomerated elements in the white and grey matters.
6 Performance of the GNN agglomeration algorithm in PDE solvers
In this section, we test the performance of the proposed agglomeration strategy when applied within PDE solvers. Specifically, we consider a three-dimensional linear elasticity equation (with unitary Lamé coefficients) with Dirichlet boundary conditions, which we discretize using a polytopal discontinuous Galerkin method. Our model problem reads:
$\begin{dcases}-\nabla\cdot(2\boldsymbol{\varepsilon}(\boldsymbol{u})+(\nabla\cdot\boldsymbol{u})\mathbf{I})=\boldsymbol{f}&\mathrm{in}\,\Omega,\\ \boldsymbol{u}=\boldsymbol{g}&\mathrm{on}\,\partial\Omega,\end{dcases}$ (10)
where $\boldsymbol{\varepsilon}(\boldsymbol{u})$ is the symmetric part of the gradient of the displacement $\boldsymbol{u}\in\mathbf{H}^{1}(\Omega)$, $\boldsymbol{f}\in\mathbf{L}^{2}(\Omega)$ is the given forcing term, and $\boldsymbol{g}\in\mathbf{H}^{1/2}(\partial\Omega)$ is the Dirichlet boundary datum. The discretization of problem (10) is based on a polytopal discontinuous Galerkin method, cf. [68] for the method’s formulation and analysis of the corresponding time-dependent problem. The discrete approximation space is defined as $\boldsymbol{V}_{h}^{\ell}=\{\boldsymbol{v}\in\mathbf{L}^{2}(\Omega):\boldsymbol{v}|_{K}\in[\mathbb{P}^{\ell}(K)]^{3}\,\forall K\in\mathscr{T}_{h}\}$, where $\ell\geq 1$ is the polynomial approximation order. In the DG formulation, we choose the stabilization parameter to be equal to $10$. The subsequent numerical tests are performed considering the following manufactured solution:
$\boldsymbol{u}(x,y,z)=\left[\cos(\pi x)\sin(\pi y)\sin(\pi z),\sin(\pi x)\cos(\pi y)\sin(\pi z),\sin(\pi x)\sin(\pi y)\cos(\pi z)\right]^{T}.$
The forcing term $\boldsymbol{f}$ and the boundary condition $\boldsymbol{g}$ are computed accordingly. We compute the discretization errors as the difference between the exact solution $\boldsymbol{u}\in\mathbf{H}^{1}(\Omega)$ and the numerical solution $\boldsymbol{u}_{h}\in\boldsymbol{V}_{h}^{\ell}$ in both the $\mathbf{L}^{2}$-norm $\|\cdot\|_{\mathbf{L}^{2}}$ and the DG-norm, defined as $\|\boldsymbol{\cdot}\|_{\mathrm{DG}}^{2}=\sum_{K\in\mathscr{T}_{h}}\|\boldsymbol{\varepsilon}(\boldsymbol{\cdot})\|_{\mathbf{L}^{2}(K)}^{2}+\sum_{F\in\mathscr{F}_{h}}\|\eta^{1/2}[\![\boldsymbol{\cdot}]\!]\|_{\mathbf{L}^{2}(F)}^{2}$, where $\![\cdot$]\!] denotes the jump operator on the skeleton $\mathscr{F}_{h}$ of the mesh, cf. [68]
6.1 Test case 6: convergence test on tetrahedral and agglomerated grids
We consider a cubic domain $\Omega=(0,1)^{3}$ discretized with a tetrahedral mesh $\mathscr{T}_{h}^{1}$ consisting of 1 3561\,356 elements (see Figure 14(a). Using our GNN agglomeration algorithm and setting as desired mesh size $0.5\textrm{diam}(\Omega)$, we construct a polyhedral mesh $\mathscr{T}_{h}^{2}$ of $39$ elements (see Figure 14(b). We solve problem (10) on the mesh $\mathscr{T}_{h}^{1}$ and varying the polynomial approximation order $\ell=1,2,3$. We repeated the same simulation on the mesh $\mathscr{T}_{h}^{2}$ with $\ell=1,...,5$. The computed approximation errors for the two simulations are reported in Figure 14(c), where the errors are plotted versus the cubic root of the total number of degrees of freedom DoFs (log-log scale). We observe that, for the considered test case—where the underlying manufactured solution is smooth—the approximation on the agglomerated mesh yields better results than that on the initial grid at a fixed level of accuracy. Moreover, this test case confirms that the proposed agglomeration strategy can also be effectively used in combination with a PDE solver.
(a) Tetrahedral mesh $\mathscr{T}_{h}^{1}$.
(b) Agglomerated mesh $\mathscr{T}_{h}^{2}$.
[…]
(c) Computed errors.
Figure 14: Test case 6: (a) initial tetrahedral mesh; (b) corresponding agglomerated mesh; (c) computed errors measured in the $\mathbf{L}^{2}$- and DG-norms versus the cubic root of the total number of degrees of freedom DOFs (loglog scale).
6.2 Test case 7: agglomeration of meshes with badly-shaped elements
The last test case considers the solution of the linear elasticity problem (10) on the bolt domain $\Omega_{2}$ considered in Section 5. The initial tetrahedral mesh, consisting of 12,175 elements, was downloaded from the FTetWild database [64] (see Figure 15(b)). We note that this mesh contains severely distorted elements, and the goal of this test is to demonstrate that the proposed GNN agglomeration algorithm can also improve overall mesh quality. Using our algorithm, we construct a polyhedral mesh with 170 elements (see Figure 15(b)). We solve problem (10) on both the initial tetrahedral mesh with $\ell=1$ and the agglomerated mesh for $\ell=1,\dots,4$. Figure 15(c) reports the deformed mesh according to the computed numerical solution for $\ell=1$. The computed results align well with what is expected for the agglomerated mesh, confirming the effectiveness of the proposed GNN algorithm when used within a PDE solver. We also compute the approximation errors in both the $\mathbf{L}^{2}$ and DG norms; the results are reported in Table 8. On the agglomerated mesh, we obtain accurate results for all tested values of $\ell$. In contrast, the initial tetrahedral mesh fails to produce a stable numerical solution due to the presence of severely distorted elements, which negatively affect the solver’s accuracy. These results suggest that the agglomeration process not only enables accurate solutions but also improves mesh quality by reducing the occurrence of highly stretched tetrahedra.
(a) Tetrahedral mesh.
(b) Agglomerated mesh.
(c) Mesh deformed by $\boldsymbol{u}_{h}$ for $\ell=1$.
Figure 15: Test case 6: (a) initial tetrahedral mesh; (b) corresponding agglomerated mesh; (c) agglomerated mesh deformed according to the computed numerical solution for $\ell=1$.
Computed errors on the agglomerated mesh
Polynomial Degree Number of DOFs $\mathbf{L}^{2}$-Error DG-Error
$\ell=1$ 2 0402\,040 $1.54\times 10^{-2}$ $7.94\times 10^{-1}$
$\ell=2$ 5 1005\,100 $4.08\times 10^{-3}$ $2.69\times 10^{-1}$
$\ell=3$ 10 20010\,200 $5.44\times 10^{-4}$ $5.74\times 10^{-2}$
$\ell=4$ 17 85017\,850 $5.85\times 10^{-5}$ $8.59\times 10^{-3}$
Computed errors on the tetrahedral mesh
Polynomial Degree Number of DOFs $\mathbf{L}^{2}$-Error DG-Error
$\ell=1$ 146 1000146\,1000 $3.44\times 10^{-1}$ $1.61\times 10^{+2}$
Table 8: Test case 7: computed approximation errors for the agglomerated mesh with $\ell=1,...,4$ and the tetrahedral mesh with $\ell=1$.
@@ 7 Conclusions
In this work, we presented a novel agglomeration algorithm for three-dimensional grids using a graph partitioning approach based on Graph Neural Networks (GNNs). Our model, which extends the results of [27], learns the geometric structure of mesh elements automatically and efficiently, while remaining independent of specific geometries or underlying differential models. This flexibility enables the method to handle complex scenarios—such as heterogeneous domains and microstructures—that classical strategies cannot process automatically, while preserving the geometric quality of the original mesh and significantly reducing computational costs. We tested the proposed approach across diverse test cases, including real geometries reconstructed from medical images and domains with microstructures and/or featuring complex geometric details. We compared the results with the analogous ones obtained by employing graph partitioners such as METIS and k-means on various test cases. Results show that our method maintains high mesh quality while efficiently handling both the geometry, the mesh topology, and the underlying physical parameters. Additionally, the GNN model offers lower online computational cost. The proposed approach has also been tested when used within PDE solvers, confirming its robustness and generalization capabilities.
Future work will focus on exploring reinforcement learning frameworks, as developed by [46], as well as extending the benchmarking considering domain-specific partitioning strategies such as SCOTCH [51, 52] or Coarsening Schemes for Graph Partitioning (e.g. [49, 50])